# A dimensional mass transference principle from balls to open sets and applications to dynamical Diophantine approximation

### v1

@@ 4.3. Application to shrinking target problems
where $s$ satisfies $P(-s(f+\log\beta),T_{\beta})=0$.
@@
Note that Lipschitz functions are dense in $C^{0}([0,1])$. Approximating $f$ from above and using the continuity of pressure function, we can assume that $f$ is Lipschitz. Let $\nu_{s}$ be the Gibbs measure associated to the Lipschitz function $s(f+\log\beta)$. The Gibbs property of $\nu_{s}$ (see Theorem 4.7) ensures that $\nu_{s}$ is quasi-self-conformal with respect to the collection of full cylinders. Moreover, by Lemma 4.4, the $\limsup$ set defined by the collection of full cylinders has full Lebesgue. Therefore, Theorem 1.4 can be applied to $\nu_{s}$.
Birkhoff’s ergodic theorem gives that for $\nu_{s}$ almost all $x$,
@@
Since for any $x\in[0,1)$, there exist infinitely many $n$ such that $I_{n,\beta}(x)$ is full, it is not difficult to verify that the set
@@
is of full $\nu_{s}$-measure, where recall that $\Lambda_{\beta}^{n}$ is the set of $n$th level full sequences, and
@@
where $E_{n}(T_{\beta},f,h_{n})=\{x\in[0,1):|T_{\beta}^{n}x-h_{n}(x)|<e^{-S_{n}f(x)}\}$. We need the following lemma to estimate the size of $I_{n,\beta}(\bm{\epsilon}_{n})\cap E_{n}(T_{\beta},f,h_{n})$.
Lemma 4.9 ([42]).
Let $h$ be a Lipschitz function with Lipschitz constant $L\geq 0$. Let $0<r<1$. For any $n$ with $L<\beta^{n}$ and any sequence $\bm{\epsilon}_{n}\in\Sigma_{\beta}^{n}$, the set
$\{x\in I_{n,\beta}(\bm{\epsilon}_{n}):|T_{\beta}^{n}x-h(x)|<r\}$
By Lipschitz continuity, for any $x,y\in I_{n,\beta}(\bm{\epsilon}_{n})$,
$|S_{n}f(x)-S_{n}f(y)|\ll 1\quad\Longrightarrow\quad e^{-S_{n}f(x)}\asymp e^{-S_{n}f(y)}.$
Therefore, since $\bm{\epsilon}_{n}\in\Lambda_{\beta}^{n}(\nu_{s},\varepsilon)$ is full, we can apply the above lemma to conclude that $I_{n,\beta}(\bm{\epsilon}_{n})\cap E_{n}(T_{\beta},f,h_{n})$ contains an interval of length
$\asymp\beta^{-n}e^{-S_{n}f(x)}\geq\beta^{-n}e^{-n\int f\,\mathrm{d}\nu_{s}-n\varepsilon}=\beta^{-n-n(\int f\,\mathrm{d}\nu_{s}+\varepsilon)/\log\beta}.$
By the variational principle and note that $P(-s(f+\log\beta),T_{\beta})=0$, we have
@@
Let $K$ be a sufficiently large integer, independent of $\varepsilon$, such that the following inequality holds:
$\bigg{(}1+\frac{\int f\,\mathrm{d}\nu_{s}+\varepsilon}{\log\beta}\bigg{)}\bigg{(}\frac{h_{\nu_{s}}}{\int f\,\mathrm{d}\nu_{s}+\log\beta}-K\varepsilon\bigg{)}\leq\frac{h_{\nu_{s}}}{\log\beta}.$
Then, by Lemma 4.8,
$\begin{split}\mathcal{H}_{\infty}^{s-K\varepsilon}\big{(}I_{n,\beta}(\bm{\epsilon}_{n})\cap E_{n}(T_{\beta},f,h_{n})\big{)}&\gg\beta^{-n(1+(\int f\,\mathrm{d}\nu_{s}+\varepsilon)/\log\beta)(s-K\varepsilon)}\\ &\geq\beta^{-nh_{\nu_{s}}/\log\beta}=|I_{n,\beta}(\bm{\epsilon}_{n})|^{-\operatorname{\dim_{H}}\nu_{s}}.\end{split}$
@@ 5.2. Pressure function
Let $\phi:[0,1]\to\mathbb{R}$ be a Lipschitz continuous function. The pressure function for Gauss map associated to $\phi$ is defined as
@@
where $y\in I_{n}(\bm{a}_{n})$. The proof of the existence of limit in (5.3) can be found in [32, Proposition 2.4]. Compared with the $\beta$-transformation, one major difference is that there are infinitely many $n$th level cylinders. As a result, the summation in (5.3) may be infinite, and hence the pressure function may fail to be continuous with respect to $\phi$. For this reason, instead of providing a comprehensive but technically involved description of the pressure function for the Gauss map, we merely summarize part of the results from [40, 34, 18] and refer the reader to these references for further details.
Theorem 5.4 ([40, 34, 18]).
Let $f:[0,1]\to\mathbb{R}$ be a positive Lipschitz function. Then the function $t\mapsto P(-t(f+\log|G^{\prime}|),G)$ is continuous on $(1/2,+\infty)$. Moreover, there exists $t=t(f)\in(1/2,1)$ such that
@@
$S=\bigg{\{}\phi_{i}(x)=\frac{1}{i+x}:i\in\mathbb{N}\bigg{\}}.$
It then follows from [34, Proposition 3.3] that the pressure function $P(-t\log|G^{\prime}|,G)$ is continuous on $(1/2,\infty)$, since by [40, Lemma 2.6], we have
$\lim_{u\to 1/2^{+}}P(-u\log|G^{\prime}|,G)=\infty.$
Now, observe that
$\begin{split}P(-t\log|G^{\prime}|,G)-t\max_{x\in[0,1]}f(x)&\leq P(-t(f+\log|G^{\prime}|),G)\\ &\leq P(-t\log|G^{\prime}|,G)-t\min_{x\in[0,1]}f(x).\end{split}$
Since $f$ is positive and $P(-\log|G^{\prime}|,G)=0$, and noting that the pressure function is strictly decreasing in $t$, there must exist a unique $t\in(1/2,1)$ such that
@@
Finally, the Lipschitz continuity of $f$ ensures the existence and uniqueness of the equilibrium state associated with the potential $t(f+\log|G^{\prime}|)$; see [18, Theorem 2.16]. In addition, this equilibrium state satisfies the corresponding Gibbs property; see [18, (2.16’)]. ∎
It follows from ergodic theorem and Lemma [17, Lemma 2.12 (b)] that for the Gibbs measure $\nu_{t}$ given in Theorem 5.4,
(5.4) $\operatorname{\dim_{H}}\nu_{t}=\frac{h_{\nu_{t}}}{\int\log|T^{\prime}|\,\mathrm{d}\nu_{t}}.$
@@ 5.3. Applications
Note that for any $\bm{a}_{n}\in\mathbb{N}^{n}$, it is known (see, e.g., [26, Lemma 2.5]) that
$G^{n}|_{I_{n}(\bm{a}_{n})}=[0,1)\mathrm{\quad and\quad}|(G^{n})^{\prime}(x)|\asymp q(\bm{a}_{n})^{2}.$
The Gibbs property further implies that $\nu_{t}$ is quasi-self-conformal with respect to the collection of all cylinders. By a suitable arrangement, it is easy to verify that the $\limsup$ set defined by the collection of cylinders has full Lebesgue measure. This key observation enables us to apply Theorem 1.4 to the measure $\nu_{t}$. In light of Theorem 5.4 and the dimension formula (5.4), we conclude — by arguments similar to those used in the proof of the large intersection property of $W(T_{\beta},f,\{h_{n}\})$ (see §4.3)— that for any continuous positive function $f:[0,1]\to\mathbb{R}$,
@@
Let us now consider the two functions appearing on the right-hand-side. On the one hand, as shown in the proof of [40, Lemma 2.6],
@@
By Birkhoff’s ergodic theorem, for $\nu_{u}$ almost all $x$,
@@
For any $\varepsilon>0$, it is not difficult to verify that the set
$\bigcap_{N=1}^{\infty}\bigcup_{n=N}^{\infty}\bigcup_{\bm{a}_{n}\in\Gamma_{n}(\nu_{u},\varepsilon)}I_{n}(\bm{a}_{n})$
@@
To analyze the growth condition on the partial quotients, define the sequence
$\alpha_{i}=B^{g_{m}(u)(1-u)^{i-1}u^{-i}}\quad\text{for $1\leq i\leq m-1$},$
and set
$\alpha_{m}=\frac{B}{\alpha_{1}\cdots\alpha_{m-1}}.$
@@
with either $a_{n+k}<a_{n+k}^{\prime\prime}<a_{n+k}^{\prime}$ or $a_{n+k}^{\prime}<a_{n+k}^{\prime\prime}<a_{n+k}$, which lies between the two cylinders stated in the lemma. Therefore, by Lemma 5.2, (5.2) and (5.8), they are separated by a distance
@@
where $\mathcal{L}$ denotes the one-dimensional Lebesgue measure.
@@
where $K$ is a sufficiently large integer, independent of $\varepsilon$, such that
$u-K\varepsilon+\frac{g_{m}(u)\log B}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}-\varepsilon}\leq u+\frac{g_{m}(u)\log B}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}}.$
@@
then trivially,
$\lambda(B(x,r))\leq 1\ll\frac{r^{u-K\varepsilon}}{|I_{n}(\bm{a}_{n})|^{u-K\varepsilon}}\ll r^{u-K\varepsilon}q(\bm{a}_{n})^{2(u-K\varepsilon)}B^{ng_{m}(u)}.$
@@
    From the well-separation property (Lemma 5.6), cylinders of level $n+k$ in $A$ are spaced by at least $1/(32\,q(\bm{a}_{n},a_{n+1},\dots,a_{n+k})^{2})$. Thus,
    (5.11) $\Delta(x;k)\ll r\,q(\bm{a}_{n},a_{n+1},\dots,a_{n+k})^{2}\asymp r\,q(\bm{a}_{n})^{2}a_{n+1}^{2}\cdots a_{n+k}^{2}.$
Combining (5.10) and (5.11) and using the inequality $\max\{a,b\}\leq a^{1-u}b^{u}$, we get
$\begin{split}\#\Delta(x;k)&\ll\max\{\alpha_{k}^{n},rq(\bm{a}_{n})^{2}a_{n+1}^{2}\cdots a_{n+k}^{2}\}\\ &\ll\alpha_{k}^{n(1-u)}\cdot\left(rq(\bm{a}_{n})^{2}a_{n+1}^{2}\cdots a_{n+k}^{2}\right)^{u}\\ &\ll r^{u}q(\bm{a}_{n})^{2u}\alpha_{1}^{2nu}\cdots\alpha_{k-1}^{2nu}\alpha_{k}^{n(1+u)},\end{split}$
where we use $a_{n+i}\asymp\alpha_{i}^{n}$ for $1\leq i\leq m$ in the last inequality. Therefore, by the definition of $\lambda$ and the equalities presented in (5.7),
$\begin{split}\lambda(B(x,r))&\ll\#\Delta(x;k)\cdot\frac{1}{\alpha_{1}^{n}\cdots\alpha_{k}^{n}}\\ &\ll r^{u}q(\bm{a}_{n})^{2u}\alpha_{1}^{n(2u-1)}\cdots\alpha_{k-1}^{n(2u-1)}\alpha_{k}^{nu}\\ &=r^{u}q(\bm{a}_{n})^{2u}B^{ng_{m}(u)}\ll r^{u-K\varepsilon}q(\bm{a}_{n})^{2(u-K\varepsilon)}B^{ng_{m}(u)}.\end{split}$
@@
$\begin{split}\frac{\mathcal{L}|_{I_{n+m}}(B(x,r))}{\mathcal{L}(I_{n+m})}&\leq\frac{2r}{\mathcal{L}(I_{n+m})}\ll rq(\bm{a}_{n},a_{n+1},\dots,a_{n+m})^{2}\\ &\ll r^{u}q(\bm{a}_{n},a_{n+1},\dots,a_{n+m})^{2u}\asymp r^{u}q(\bm{a}_{n})^{2u}\alpha_{1}^{2nu}\cdots\alpha_{m}^{2nu}\\ &\leq r^{u}q(\bm{a}_{n})^{2u}\alpha_{1}^{2nu}\cdots\alpha_{m-1}^{2nu}\alpha_{m}^{n(u+1)},\end{split}$
where we use $1/2<u<1$ in the last inequality. Again, by the definition of $\lambda$ and the equalities presented in (5.7),
$\begin{split}\lambda(B(x,r))&=\frac{1}{\#A}\cdot\frac{\mathcal{L}|_{I_{n+m}}(B(x,r))}{\mathcal{L}(I_{n+m})}\ll r^{u}q(\bm{a}_{n})^{2u}\alpha_{1}^{n(2u-1)}\cdots\alpha_{m-1}^{n(2u-1)}\alpha_{m}^{nu}\\ &=r^{u}q(\bm{a}_{n})^{2u}B^{ng_{m}(u)}\ll r^{u-K\varepsilon}q(\bm{a}_{n})^{2(u-K\varepsilon)}B^{ng_{m}(u)}.\qed\end{split}$
@@
$\bigg{|}\frac{1}{n}S_{n}\log|G^{\prime}|(x)-\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}\bigg{|}<\varepsilon\quad\Longrightarrow\quad S_{n}\log|G^{\prime}|(x)\geq n\bigg{(}\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}-\varepsilon\bigg{)}.$
@@
$\leq$ $S_{n}\log|G^{\prime}|(x)\bigg{(}u-K\varepsilon+\frac{ng_{m}(u)\log B}{n(\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}-\varepsilon)}\bigg{)}$
(5.13) $\leq$ $S_{n}\log|G^{\prime}|(x)\bigg{(}u+\frac{g_{m}(u)\log B}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}}\bigg{)},$
where the final inequality follows from the choice of the constant $K$. By the variational principle, we have
@@
$\mathcal{H}_{\infty}^{u-K\varepsilon}(A)\gg e^{-S_{n}\log|G^{\prime}|(x)\cdot\operatorname{\dim_{H}}\nu_{u}}\asymp|I_{n}(\bm{a}_{n})|^{\operatorname{\dim_{H}}\nu_{u}}.$
@@
$A\subset\left\{x\in I_{n}(\bm{a}_{n}):a_{n+1}(x)\cdots a_{n+m}(x)\geq B^{n}\right\},$

### v2

@@ 4.3. Application to shrinking target problems
where $s$ satisfies $P(-s(f+\log\beta),T_{\beta})=0$. The proof relies on the following simple but useful fact.
Lemma 4.12 ([47]).
Let $h:[0,1)\to[0,1)$ be a Lipschitz function with Lipschitz constant $L\geq 0$. Let $0<r<1$. For any $n$ with $L<\beta^{n}$ and any sequence $\bm{\epsilon}_{n}\in\Sigma_{\beta}^{n}$, the set
$\{x\in I_{n,\beta}(\bm{\epsilon}_{n}):|T_{\beta}^{n}x-h(x)|<r\}$
Let us turn to the proof of (4.6).
@@
Note that Lipschitz functions are dense in $C^{0}([0,1])$ and that
$W(T_{\beta},g,\{h_{n}\})\subset W(T_{\beta},f,\{h_{n}\})\quad\text{whenever $g\geq f$}.$
Since the pressure function $P(\cdot,T_{\beta})$ is continuous (see Proposition 4.9), we can choose a Lipschitz function $g$ so that the Hausdorff dimension of $W(T_{\beta},g,\{h_{n}\})$ is as close as we want to that of $W(T_{\beta},f,\{h_{n}\})$, provided (4.6) holds. Therefore, in what follows, we may assume that $f$ is Lipschitz.
Let $\nu_{s}$ be the Gibbs measure associated to the Lipschitz function $s(f+\log\beta)$. Since the concatenation of any two full sequences is still full (see Proposition 4.5), by the Gibbs property of $\nu_{s}$ (see Theorem 4.10), it is not difficult to verify that $\nu_{s}$ is quasi-self-conformal with respect to the collection of full cylinders. Moreover, by Lemma 4.7, the $\limsup$ set defined by the collection of full cylinders has full Lebesgue measure. Therefore, Theorem 1.4 can be applied to $\nu_{s}$.
Birkhoff’s ergodic theorem gives that for $\nu_{s}$-almost all $x$,
@@
Since for any $x\in[0,1)$, there exist infinitely many $n$ such that $I_{n,\beta}(x)$ is full, we have that the set
@@
is of full $\nu_{s}$-measure, where recall that $\Lambda_{\beta}^{n}$ is the set of full sequences of length $n$, and
@@
where $E_{n}(T_{\beta},f,h_{n})=\{x\in[0,1):|T_{\beta}^{n}x-h_{n}(x)|<e^{-S_{n}f(x)}\}$. By the Lipschitz continuity of $f$, for any $x,y\in I_{n,\beta}(\bm{\epsilon}_{n})$,
$|S_{n}f(x)-S_{n}f(y)|\ll 1\quad\Longrightarrow\quad e^{-S_{n}f(x)}\asymp e^{-S_{n}f(y)},$
where the implied constant is absolute. Therefore, since $\bm{\epsilon}_{n}\in\Lambda_{\beta}^{n}(\nu_{s},\varepsilon)$ is full, we can apply Lemma 4.12 to conclude that $I_{n,\beta}(\bm{\epsilon}_{n})\cap E_{n}(T_{\beta},f,h_{n})$ contains an interval of length
(4.7) $\asymp\beta^{-n}e^{-S_{n}f(x)}\geq\beta^{-n}e^{-n\int f\,\mathrm{d}\nu_{s}-n\varepsilon}=\beta^{-n(1+(\int f\,\mathrm{d}\nu_{s}+\varepsilon)/\log\beta)}.$
By the variational principle (see Theorem 4.10 (1)) and note that $P(-s(f+\log\beta),T_{\beta})=0$, we have
@@
Let $K=\frac{h_{\nu_{s}}}{(\int f\,\mathrm{d}\nu_{s}+\log\beta)\log\beta}$. Then,
$\bigg(1+\frac{\int f\,\mathrm{d}\nu_{s}+\varepsilon}{\log\beta}\bigg)\bigg(\frac{h_{\nu_{s}}}{\int f\,\mathrm{d}\nu_{s}+\log\beta}-K\varepsilon\bigg)$
$=$ $\frac{h_{\nu_{s}}}{\log\beta}+\frac{\varepsilon h_{\nu_{s}}}{(\int f\,\mathrm{d}\nu_{s}+\log\beta)\log\beta}-\bigg(1+\frac{\int f\,\mathrm{d}\nu_{s}+\varepsilon}{\log\beta}\bigg)\cdot K\varepsilon$
(4.8) $\leq$ $\frac{h_{\nu_{s}}}{\log\beta}+\frac{\varepsilon h_{\nu_{s}}}{(\int f\,\mathrm{d}\nu_{s}+\log\beta)\log\beta}-K\varepsilon=\frac{h_{v_{s}}}{\log\beta}.$
This together with Lemma 4.11 yields
[…]
@@
Remark 10.
In order to highlight further applications, such as to expanding Markov maps with finite partitions on $[0,1]$, of Theorem 1.4, we summarize the properties used in the proof of (4.6) as follows:
  1. (1)
    There exists $s\geq 0$ such that $P(-s(f+\log\beta),T_{\beta})=0$.
  2. (2)
    For the pressure function $s(f+\log\beta)$ given in item (1), there exists a Gibbs measure $\nu_{s}$ verifies the variational principle
    (4.9) $0=P(-s(f+\log\beta),T_{\beta})=h_{\nu_{s}}+\int-s(f+\log\beta)\,d\nu_{s}.$
    The Gibbs property of $\nu_{s}$ ensures that this measure is quasi-self-conformal with respect to the collection of all cylinders.
  3. (3)
    The dimension of $\nu_{s}$ satisfies
[…]
    where $T_{\beta}^{\prime}$ denotes the derivative of $T_{\beta}$.
These results are quite standard in the study of multifractal analysis, especially for expanding Markov maps with finite partitions on $[0,1]$. For a summary, see, for example, [35, §2.3]. Although the content described in that reference differs from our paper, it does indeed lead to items (1)–(3), with $T_{\beta}$ and $\log\beta$ replaced by the Markov map $T$ and the logarithm of its derivative $\log|T^{\prime}|$, respectively. To keep the paper at a manageable length, we leave the verification of these details to the interested reader.
@@ 5.1. Definition and some basic properties
By the recursive formulae (5.2), for any $(a_{1},\dots,a_{n})\in\mathbb{N}^{n}$,
(5.3) $q(a_{1},\dots,a_{n-1})a_{n}\leq q(a_{1},\dots,a_{n})\leq q(a_{1},\dots,a_{n-1})(a_{n}+1).$
@@ 5.2. Pressure function
For any $n\in\mathbb{N}$ and any function $\phi:[0,1)\to\mathbb{R}$, the $n$th variation of $\phi$ is defined by
$\operatorname{Var}_{n}(\phi):=\sup\{|\phi(x)-\phi(y)|:I_{n}(x)=I_{n}(y)\},$
where $I_{n}(x)$ denotes the $n$th level cylinder containing $x$. Let $\phi:[0,1)\to\mathbb{R}$ be a real function (not necessarily continuous) with $\operatorname{Var}_{1}(\phi)<\infty$ and $\operatorname{Var}_{n}(\phi)\to 0$ as $n\to\infty$. The pressure function for Gauss map associated to $\phi$ is defined as
@@
where $y\in I_{n}(\bm{a}_{n})$. The proof of the existence of limit in (5.4) can be found in [34, Proposition 2.4]. Clearly, if $\phi$ is either $\log|G^{\prime}|$ or a continuous function on $[0,1]$, then $\operatorname{Var}_{1}(\phi)<\infty$ and $\operatorname{Var}_{n}(\phi)\to 0$ as $n\to\infty$. Therefore, the pressure for such $\phi$ exists.
Compared with the $\beta$-transformations, one major difference is that there are infinitely many $n$th level cylinders. As a result, the summation in (5.4) may be infinite, and hence the pressure function may fail to be continuous with respect to $\phi$. For this reason, instead of providing a comprehensive but technically involved description of the pressure function for the Gauss map, we merely summarize part of the results from [45, 37, 20] and refer the reader to these references for further details.
Lemma 5.4 ([37, Proposition 3.3] and [45, Lemma 2.6]).
The function $t\mapsto P(-t\log|G^{\prime}|,G)$ is strictly decreasing, convex and continuous on $(1/2,\infty)$ and satisfies
$\lim_{t\to 1/2^{+}}P(-t\log|G^{\prime}|,G)=\infty\mathrm{\quad and\quad}P(-\log|G^{\prime}|,G)=0.$
Proof.
$S=\bigg\{\phi_{i}(x)=\frac{1}{i+x}:[0,1]\to[0,1]:i\in\mathbb{N}\bigg\}.$
It then follows from [37, Proposition 3.3] that the pressure function $P(-t\log|G^{\prime}|,G)$ is strictly decreasing, convex and continuous on $(\theta_{S},\infty)$ with
$\theta_{S}:=\inf\{t:P(-t\log|G^{\prime}|,G)<\infty\}.$
For any $n\in\mathbb{N}$, by Lemma 5.2
$\sum_{\bm{a}_{n}\in\mathbb{N}^{n}}e^{-tS_{n}\log|G^{\prime}|(y)}\asymp\sum_{\bm{a}_{n}\in\mathbb{N}^{n}}|I_{n}(\bm{a}_{n})|^{t}.$
Hence, if $t=1$, then by the definition of cylinders, the above sum is approximately $1$, and so $P(-\log|G^{\prime}|,G)=0$. Now, suppose that $t<1$. Then, by Lemma 5.2 and (5.3),
$\sum_{\bm{a}_{n}\in\mathbb{N}^{n}}|I_{n}(\bm{a}_{n})|^{t}\asymp\sum_{\bm{a}_{n}\in\mathbb{N}^{n}}q(\bm{a}_{n})^{-2t}\leq\sum_{\bm{a}_{n}\in\mathbb{N}^{n}}\bigg(\frac{1}{a_{1}\cdots a_{n}}\bigg)^{2t}=\bigg(\sum_{k=1}^{\infty}\frac{1}{k^{2t}}\bigg)^{n}$
and similarly,
$\sum_{\bm{a}_{n}\in\mathbb{N}^{n}}|I_{n}(\bm{a}_{n})|^{t}\gg\bigg(\sum_{k=1}^{\infty}\frac{1}{(k+1)^{2t}}\bigg)^{n}.$
This immediately implies that $P(-t\log|G^{\prime}|,G)<\infty$ if and only if $t>1/2$. Therefore, $\theta_{S}=1/2$ and the lemma follows. ∎
Theorem 5.5 ([37, 20]).
Let $f:[0,1]\to\mathbb{R}$ be a positive Lipschitz continuous function. Then the function $t\mapsto P(-t(f+\log|G^{\prime}|),G)$ is continuous on $(1/2,+\infty)$. Moreover, the following statements hold:
  1. (1)
    there exists $t=t(f)\in(1/2,1)$ such that
@@
  3. (3)
    the Hausdorff dimension of $\nu_{t}$ satisfies
    (5.6) $\operatorname{\dim_{H}}\nu_{t}=\frac{h_{\nu_{t}}}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{t}}.$
@@
(1) Observe that
$\begin{split}P(-t\log|G^{\prime}|,G)-t\|f\|_{\infty}&\leq P(-t(f+\log|G^{\prime}|),G)\\ &\leq P(-t\log|G^{\prime}|,G)-t\min_{x\in[0,1]}f(x).\end{split}$
By Lemma 5.4, we have
$\lim_{t\to 1/2^{+}}P(-t(f+\log|G^{\prime}|),G)\geq\lim_{t\to 1/2^{+}}P(-t\log|G^{\prime}|,G)-t\|f\|_{\infty}=\infty$
and
$P(-(f+\log|G^{\prime}|),G)\leq P(-\log|G^{\prime}|,G)-\min_{x\in[0,1]}f(x)\leq 0.$
Since the pressure function is strictly decreasing and continuous in $t$ (see Lemma 5.4 again), there must exist a unique $t\in(1/2,1)$ such that
@@
(2) The Lipschitz continuity and positivity of $f$ together ensures the existence and uniqueness of the equilibrium state associated with the potential $t(f+\log|G^{\prime}|)$; see [20, Theorem 2.16]. The variational principle (5.5) follows from their definition of the equilibrium state (see [20, (2.19)]) directly. In addition, this equilibrium state satisfies the corresponding Gibbs property; see [20, (2.16’)].
(3) This is a consequence of Birkhoff’s ergodic theorem and [18, Lemma 2.12 (b)]. ∎
@@ 5.3. Applications
Note that for any $\bm{a}_{n}\in\mathbb{N}^{n}$, it is known (see, e.g., [28, Lemma 2.5]) that
$G^{n}|_{I_{n}(\bm{a}_{n})}=[0,1)\mathrm{\quad and\quad}|(G^{n})^{\prime}(x)|\asymp q(\bm{a}_{n})^{2},$
where $x\in I_{n}(\bm{a}_{n})$. The Gibbs property (see Theorem 5.5) further implies that $\nu_{t}$ is quasi-self-conformal with respect to the collection of all cylinders. By a suitable arrangement, it is easy to verify that the $\limsup$ set defined by the collection of cylinders has full Lebesgue measure. This key observation enables us to apply Theorem 1.4 to the measure $\nu_{t}$. In light of Theorem 5.5, we conclude — by arguments similar to those used in the proof of the large intersection property of $W(T_{\beta},f,\{h_{n}\})$ (see §4.3 and Remark 10)— that for any positive continuous function $f:[0,1]\to\mathbb{R}$,
@@
Let us now consider the two functions of $u$ appearing on the right-hand-side. On the one hand, by Lemma 5.4,
@@
By Birkhoff’s ergodic theorem, for $\nu_{u}$-almost all $x$,
@@
For any $0<\varepsilon<\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}/2$, it is not difficult to verify that the set
(5.9) $\bigcap_{N=1}^{\infty}\bigcup_{n=N}^{\infty}\bigcup_{\bm{a}_{n}\in\Gamma_{n}(\nu_{u},\varepsilon)}I_{n}(\bm{a}_{n})$
@@
Let
$\alpha_{i}=B^{g_{m}(u)(1-u)^{i-1}u^{-i}}\quad\text{for $1\leq i\leq m-1$},\quad\text{and}\quad\alpha_{m}=\frac{B}{\alpha_{1}\cdots\alpha_{m-1}}.$
@@
with either $a_{n+k}<a_{n+k}^{\prime\prime}<a_{n+k}^{\prime}$ or $a_{n+k}^{\prime}<a_{n+k}^{\prime\prime}<a_{n+k}$, lies between the two cylinders stated in the lemma. Therefore, by Lemma 5.2, (5.2) and (5.11), they are separated by a distance
@@
where $\mathcal{L}$ denotes the one-dimensional Lebesgue measure. That is, we assign each $(n+m)$th level cylinder of equal weight. For $1\leq k\leq m$, the number of descendants of each $(\bm{a}_{n},a_{n+1},\dots,a_{n+k})$ are the same. Therefore, we have
(5.13) $\lambda(I_{n+k}(\bm{a}_{n},a_{n+1},\dots,a_{n+k}))\asymp\frac{1}{\alpha_{1}^{n}\cdots\alpha_{m}^{n}}\cdot\alpha_{k+1}^{n}\cdots\alpha_{m}^{n}=\frac{1}{\alpha_{1}^{n}\cdots\alpha_{k}^{n}}.$
@@
where $K=\frac{g_{m}(u)\log B}{2(\int\log|G^{\prime}|\,\mathrm{d}\nu_{u})^{2}}$.
@@
then by Lemma 5.2,
$\lambda(B(x,r))\leq 1\ll\frac{r^{u-K\varepsilon}}{|I_{n}(\bm{a}_{n})|^{u-K\varepsilon}}\asymp r^{u-K\varepsilon}q(\bm{a}_{n})^{2(u-K\varepsilon)}\leq r^{u-K\varepsilon}q(\bm{a}_{n})^{2(u-K\varepsilon)}B^{ng_{m}(u)}.$
@@
    From the well-separation property (see Lemma 5.7), cylinders of level $n+k$ in $A$ are spaced by at least $1/(32\,q(\bm{a}_{n},a_{n+1},\dots,a_{n+k})^{2})$. Thus,
[…]
Combining (5.14) and (5.15) and using the inequality $\min\{a,b\}\leq a^{1-u}b^{u}$, we get
$\#\Delta(x;k)$ $\ll\min\{\alpha_{k}^{n},rq(\bm{a}_{n})^{2}a_{n+1}^{2}\cdots a_{n+k}^{2}\}$
$\ll\alpha_{k}^{n(1-u)}\cdot\left(rq(\bm{a}_{n})^{2}a_{n+1}^{2}\cdots a_{n+k}^{2}\right)^{u}$
(5.16) $\ll r^{u}q(\bm{a}_{n})^{2u}\alpha_{1}^{2nu}\cdots\alpha_{k-1}^{2nu}\alpha_{k}^{n(1+u)},$
where we use $a_{n+i}\asymp\alpha_{i}^{n}$ for $1\leq i\leq m$ in the last inequality. Hence,
[…]
where the last inequality follows from $r\ll q(\bm{a}_{n})$.
@@
[…]
where we use $1/2<u<1$ in the last inequality. This together with the definition of $\lambda$ gives
[…]
@@
$\bigg|\frac{1}{n}S_{n}\log|G^{\prime}|(x)-\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}\bigg|<\varepsilon~\Longrightarrow~S_{n}\log|G^{\prime}|(x)\geq n\bigg(\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}-\varepsilon\bigg).$
@@
$\leq$ $S_{n}\log|G^{\prime}|(x)\bigg(u-K\varepsilon+\frac{ng_{m}(u)\log B}{n(\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}-\varepsilon)}\bigg).$
Recall that $0<\varepsilon<\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}/2$ (see the line before (5.9)) and $K=\frac{g_{m}(u)\log B}{2(\int\log|G^{\prime}|\,\mathrm{d}\nu_{u})^{2}}$. It follows that
$\begin{split}&\frac{ng_{m}(u)\log B}{n(\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}-\varepsilon)}-K\varepsilon\\ =&\frac{g_{m}(u)\log B}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}}\cdot\frac{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}-\varepsilon}-\frac{g_{m}(u)\log B}{2(\int\log|G^{\prime}|\,\mathrm{d}\nu_{u})^{2}}\cdot\varepsilon\\ =&\frac{g_{m}(u)\log B}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}}+\frac{g_{m}(u)\log B}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}}\cdot\frac{\varepsilon}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}-\varepsilon}-\frac{g_{m}(u)\log B}{2(\int\log|G^{\prime}|\,\mathrm{d}\nu_{u})^{2}}\cdot\varepsilon\\ \leq&\frac{g_{m}(u)\log B}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}}.\end{split}$
Therefore,
(5.18) $(u-K\varepsilon)S_{n}\log|G^{\prime}|(x)+ng_{m}(u)\log B\leq S_{n}\log|G^{\prime}|(x)\bigg(u+\frac{g_{m}(u)\log B}{\int\log|G^{\prime}|\,\mathrm{d}\nu_{u}}\bigg).$
By the variational principle (see Theorem 5.5 (2)), we have
@@
[…]
@@
$A\subset\left\{x\in I_{n}(\bm{a}_{n}):a_{n+1}(x)\cdots a_{n+m}(x)\geq B^{n}\right\}\mathrm{\quad and\quad}\bm{a}_{n}\in\Gamma_{n}(\nu_{u},\varepsilon),$